How many moles of hydrogen are in 2.0 mole(s) of quinin? Express your answer using two significant figures?

How many moles of carbon are in 4.6 moles of quinine?
Express your answer using two significant figures.

How many moles of nitrogen are in 1.5×10−2 mole of quinine?
Express your answer using two significant figures.
1
Expert's answer
2018-11-05T09:52:09-0500

1. How many moles of hydrogen are in 2.0 mole(s) of quinin? Express your answer using two significant figures?

Solution.

Formula: C20H24N2O2

n(H) = n(C20H24N2O2)*N(H) = 2.0*24 = 48 moles

Answer: 48 moles

2. How many moles of carbon are in 4.6 moles of quinine? Express your answer using two significant figures.

Solution.

Formula: C20H24N2O2

n(C) = n(C20H24N2O2)*N(C) = 4.6*20 = 92 moles

Answer: 92 moles

3. How many moles of nitrogen are in 1.5×10−2 mole of quinine? Express your answer using two significant figures.

Solution.

Formula: C20H24N2O2

n(N) = n(C20H24N2O2)*N(N) = 0.015*2 = 0.030 moles

Answer: 0.030 moles